OpenStudy (anonymous):

If the given polynomial has a special factorization, factor the polynomial: -12y^2+36xy-27x^2

OpenStudy (anonymous):

-3(4y^2-12xy+9x^2) =-3(2y-3)^2

OpenStudy (anonymous):

factor out a -3 then get a perfect square

OpenStudy (anonymous):

\[-3(4y^2-12xy+9x^2)=-3(2y-3x)^2\]